OpenStudy (anonymous):

4x^2-3x-10=0 solve each equation by completing the square

Parth (parthkohli):

\[\textbf{HINT}\]Add \(10\) to both sides, then add \(2.25\)...

OpenStudy (anonymous):

2.25? 4x^2-3x=10

OpenStudy (anonymous):

where would i add 2.25

OpenStudy (anonymous):

This can be solved the same way I showed you before. Follow the example from your last complete the square problem.

Parth (parthkohli):

To both sides.

Parth (parthkohli):

\[4x^2 - 3x + 2.25 = 12.25\]Multiplying both sides by \(100\) is an idea.

OpenStudy (anonymous):

but where did you get 2.25 from?

Parth (parthkohli):

When you are completing the square, you basically add \(\left(\dfrac{\text{coefficient of }x}{2}\right)^2\) to both sides.

OpenStudy (anonymous):

jeez i have no idea what that means

OpenStudy (anonymous):